# Creative and geometric times in physics, mathematics, logic, and philosophy

### v1

@@ Abstract
We propose a distinction between two different concepts of time that play a role in physics: geometric time and creative time. The former is the time of deterministic physics and merely parametrizes a given evolution. The latter is instead characterized by real change, i.e. novel information that gets created when a non-necessary event becomes determined in a fundamentally indeterministic physics. This allows us to give a naturalistic characterization of the present as the moment that separates the potential future from the determined past. We discuss how these two concepts find natural applications in classical and intuitionistic mathematics, respectively, and in classical and multivalued tensed logic, as well as how they relate to the well-known A- and B-theories in the philosophy of time.
@@ I Introduction
There is almost nothing that we perceive so ubiquitously than the passage of time. And yet our most successful physical theories still struggle to make sense of this concept in an unequivocal way. Actually, modern physics has relegated time to play a less and less special role [1]. However, in the words of I. Prigogine, ``````` `no formulation of the laws of nature that does not take into account this constructive role of time can ever be satisfactory” [2].
To address this fundamental problem, we propose here two different concepts of time, which we call geometric time and creative time, respectively. We show that those stem from our fundamental assumption of physics as being either deterministic or indeterministic at the fundamental level, and that they both seem to contribute to our understanding of natural phenomena. In particular, we will see that geometric time is the parametric time appearing in the (deterministic) equations of motions of physics, just a coordinate that, together with the three spatial directions, labels fix events in a geometrical block-universe. On the other hand, creative time originates from the assumption that certain events which are fundamentally indeterminate become determinate. Our general approach is admittedly highly influenced by quantum mechanics, which has brought the concept of fundamental indeterminacy into the domain of physics. But we abstractly conceive the possibility that there is ontic indeterminacy in the world (also at the classical level, see [3, 4]). This idea of creative time resonates to some extent with the positions of the philosopher H. Bergson, well-known for having confronted Einstein about the nature of time, who stated: ``````` `Time is what prevents everything from being given all at once. […] It must therefore be development. Wouldn’t it then be the vehicle of creation and choice? Would time’s existence not prove that things are undetermined? Would time not be this very indeterminacy?” [5]. We somewhat twist this view around by assuming that it is indeterminacy that defines a different fundamental type of time. Indeed, if there is ontic indeterminacy in nature (i.e., physical pure states do not fully determine the future), this leads to multiple potential future states, i.e., to indeterminism and an open future. It is the actualization of these potentialities, i.e., when new information previously not existing is created, that makes this creative time tick.
@@
Figure 1: Typical evolution of a classical chaotic system in phase space. The initial state (a) has some ontic indeterminacy, i.e., the position and momentum are in this case not determined further than the finite region depicted in figure. When evolved through the deterministic equations of motions (b), the indeterminacy spreads (while preserving the same volume), leading to multiple potential future evolutions, i.e. to indeterminism. However, only one of these potential evolutions will actualize (c), while ruling out the other previously potential futures (dashed); it is this process of actualization that makes creative time process.
@@
Starting from physics (Sect. II), we apply these two concepts of time to different disciplines, showing that this distinction gives rise to natural parallels in mathematics (classical versus intuitionistic; Sect. III), helps clarify the logic of scientific propositions (Sect. IV), and it contributes further to the metaphysics of time (so-called A-theory versus B-theory; Sect. V). More specifically, the approach based on consecutive actualizations of potentialities, that is when creative time is at work, leads to a naturalistic characterization of the difference between past, present and future.
@@ II Two times in physics
In order to illustrate different concepts of time in physics, let us start from an example. Let us ask the following scientific question: what will be the weather like, say on the field hockey pitch of Geneva, exactly in one month from now? One can envision several procedures to reply to this question. On the one hand, it is possible to use the equations of motion of all the involved molecules of air, water vapour, etc., to calculate their evolution and so predict the weather in Geneva in a month time. On the other hand, one can wait exactly one month, let the process develop, and finally observe the weather in Geneva. In-between, one could spend one week to collect more data – or merely waste a week – and only then start computing the evolution of all air molecules. Let us concentrate on the two first mentioned extremal procedures. The question that we ask here is: Are these two procedures talking about the same unique concept of time, or are there multiple ones that play a role in physics? We contend that there are (at least) two different concepts of time, as exemplified by the two procedures above.
@@
Before elaborating on that, we ought to distinguish between different fundamental interpretations of physics, which become especially relevant in dealing with complex systems like the one determining the weather in our example. The standard narrative (i) has it that classical physics is deterministic, i.e. that each state is mapped one-to-one to all other possible states at all times (with time being the parameter that appears in the equations of motion). Although usually not stressed enough, this is based on two independent ingredients: the first is that since the equations of motion are ordinary differential equations, their solutions generally have a unique continuation in the parameter time, [^1] In exceptional cases, classical equations of motion of even simple systems do not have a unique (deterministic) solution, such as in the example of Norton’s dome [6]. However, in highly complex systems – such as Lagrangian fluid particles in high Reynolds-number turbulence – the fact that deterministic trajectories are not unique is rather common (see [7] and references therein). given appropriate initial conditions; the second is that the initial conditions exist fully determinate with infinite precision (i.e., a state is represented by a dimensionless mathematical point in phase space, in turn described by an n-tuple of real numbers). Only taking these two assumptions together justifies the purported fundamental determinism of classical physics. Another possibility, (ii) is to keep the idea that the system is fundamentally deterministic, but realistically it is not possible to know the initial conditions with infinite precision. This becomes most relevant in chaotic systems (like the weather in our example) in which the future behavior of the considered system highly depends on the infinitesimal digits of the initial conditions. In this case one can only statistically predict what are the possible future states compatible with the partial knowledge of the initial conditions. Finally, (iii) one can instead uphold the more radical assumption that the initial conditions (alternatively, the state of a system) simply do not exist determined with infinite precision, but there is an ontic indeterminacy in nature (for a collection of works on this see, [8, 3, 9, 10, 11, 4]; see also [12] and references therein for recent developments on metaphysical indeterminacy). In this case, although the equations of motion may remain unchanged, the system does not undergo a necessary evolution, but multiple future states are possible (perhaps with a determined objective tendency or propensity; see [4]), i.e., it is the ontic indeterminacy that leads to fundamental indeterminism (see Fig. 1). In what follows, when referring to indeterminacy we will have this kind of ontic property in mind and its consequent indeterminism.
@@
Let us now go back to analyze the concepts of time that are at work in the example of the weather in Geneva in a month. The first procedure, under the ideal assumption of determinism (i), merely makes use of time as a parameter that simply indicates the coordinate of a certain evolution (e.g., sunny weather in one month from the start of the computation). Together with the three spatial coordinates (i.e., in Geneva in our example), this kind of time forms a manifold that contains all the events: a four-dimensional block universe picture where past, present and future are all on the same metaphysical footing (this is also known as Eternalism in the philosophy of time; see Sect. V). This can thus be called geometric time, the time that most physicists have come to think of as fundamental, both from classical physics and even more from relativity theory where time becomes “just” another dimension like space. In this sense, time is a relational label that allows one to position an event (with respect to another point of reference), but there is no real ``````` `becoming”.
@@
On the other hand, let us consider the case of ontic indeterminacy in the physical state (iii). There, the actual development of a non-necessary physical process – such as the evolution of a highly chaotic system like the one that determines the weather, or of a quantum system – involves, in our opinion, a time that really ``````` `processes”. That is, a concept of time which sets an actual separation of the future from the past (with the present at their interface). And it does so when new information that was not existing before gets created – i.e., even having access to the whole existing information of the initial conditions (even of the whole state of the universe) before, would not determine all the future physical properties of a system afterwards. Again, we stress that we are not talking here about epistemic uncertainty, but rather we explicitly take a physicalist stance on information: it quantifies something objectively encoded in the degrees of freedom of the universe. In this sense information gets created when non-necessary physical events come about, independently of any observer. Note that, – and this is what happens also in quantum physics – although the state of the system is fully determined and it contains the maximal information about the systems (i.e., it is pure), this does not fully determine the future evolution (see also Ref. [13]). The concept of time that stems from this can thus be called creative time. [^2] The distinction between geometric and creative time was sketched in Refs. [14, 15, 3, 16] and previously also in Ref. [17] where the creative time was called ``````` `historical time”. On a similar note, G. Ellis remarked that ``````` `The most important property of time is that it unfolds. […] The time that is the present at this instant will be the past at the next instant. […] It is this fundamental feature of time that is not encapsulated by today’s theoretical physics” [18] (see also [19]).
@@
Let us emphasize that the concept of creative time has nothing to do with language, nor with subjective or psychological processes; it ticks due to objective processes that happen by themselves in nature. Accordingly, there are two types of events: necessary events (e.g., when two geometric worldlines cross each other) and non-necessary events when potentialities actualize. Those instances of actualization determine the objective present. In other words, this fundamental “becoming” comes from the realization of one single outcome among the possible ones (with the contextual exclusion of the other previously valid futures) due to ontic indeterminacy.
In the literature one often finds that time ``````` `unfolds” or that time ``````` `flows” to indicate that time is associated with something that really changes (so with creative time). Here we prefer to use the terms time ``````` `develops” or ``````` `processes”, because also a film already shot does unfold, but nothing new is created there; whereas to flow is a somewhat misleading term for it reminds of the concept of speed which has time already in it. So, henceforth, we will write that geometric time passes while creative time processes or happens.
@@
Coming back to the example of the weather in Geneva in a month, to carry out either of the envisioned procedures to determine the outcome takes some time. In the first procedure, i.e. when geometric time is at work, one maps the time that it takes to perform the computation $\Delta t^{\prime}$ to the time appearing in the equations of motion. That is, to the interval $\Delta t=t_{f}-t_{i}$, where $t_{i}$ is the parametric time associated to the initial state and $t_{f}$ is the parametric time of the considered final state (i.e., the moment when we want to know the weather in Geneva). The (deterministic) computation lasts $\Delta t^{\prime}<\Delta t=1$ month, and returns as an outcome the weather in Geneva at $t_{f}$, say ``````` `sunny”. [^3] In principle the time for the “prediction” could take longer than the time interval to observe the outcome. In such a case, one would anyways ask for consistency of the calculation afterwards.
@@
One can also think of more subtle questions about the weather, such as ``````` `will there be (anytime in history) 700 consecutive days of rain in Geneva?”. If the complex system that forms the weather is fully deterministic, answering this question seems only a matter of geometric time, although the answer may remain undecided for any finite value of geometric time. One can run the computation of the dynamics of the system and at some point find that the answer is ``````` `yes”, if one finds the 700 consecutive days of rain, otherwise the algorithm will keep searching without ever halting (however, this question might be more complex, see Sect. III).
@@
Two remarks are here in order. Firstly, note that $\Delta t^{\prime}$ can be speed up by a more efficient computation, namely, this depends on how fast one ``````` `runs” through the parameter time as it appears in the equations of motion (this depends on the used algorithm and, physically, one can really think of the speed of the head of a Turing machine in reading the tape). Moreover, time dilation in relativity can be seen as a physical process to modify the interval of geometric time (see Subsect. VI.1). So the map $\Delta t\rightarrow\Delta t^{\prime}$ is a one-to-many function. Secondly, once one reaches the final state with the associated outcome ``````` `sunny”, one gains full information about the initial state at $t_{i}$ and its consequences, but not necessarily the full information about the final state if there is some fundamental indeterminacy, i.e., the actual weather that will be observed may vary.
@@
On the contrary, the procedure involving creative time, namely, of waiting a month and carrying out an observation of the weather gives full information about the outcome, but provides in general little information about the initial state. This is because new information has been created along the way, through a series of indeterministic processes. So, the initial and the final states are not deterministically mapped to each other. The exactly same initial state under exactly the same physical conditions leads in general to different outcomes, so one can only infer a set of initial states that are compatible with the observed outcome. Creative time is thus associated to the possibility that physical properties are in general ontologically indeterminate (e.g., it is neither true that ``````` `the particle is within region $r$” nor that ``````` `the particle is not within $r$”, where $r$ is a region of space, see also Sect. IV). Creative time is required when an actualization event of the values of physical variables that were previously indeterminate occurs. It is the process of change from indeterminate to determinate – with an associated creation of novel information – that brings about this “happening” of time: creative time ticks when new information is created.

### v2

@@ Abstract
We distinguish two different concepts of time that play a role in physics: geometric time and creative time. The former is the time of deterministic physics and merely parametrizes a given evolution. The latter is instead characterized by real change, i.e. novel information that gets created when a non-necessary event becomes determinate in a fundamentally indeterministic physics. This allows one to give a naturalistic characterization of the present as the moment that separates the potential future from the determinate past. We discuss how these two concepts find natural applications in classical and intuitionistic mathematics, respectively, and in classical and intuitionistic logic, as well as how they relate to the well-known A- and B-theories in the philosophy of time. We acknowledge that we do not offer here a unified concept or a new philosophy of time. However, we contend that none of the existing philosophical accounts fully integrate both the geometric and creative concepts of time.
@@ I Introduction
There is almost nothing that we perceive so ubiquitously as the occurrence of time. And yet our most successful physical theories still struggle to make sense of this concept in an unequivocal way. Actually, modern physics has relegated time to play a less and less special role [82]. However, in the words of I. Prigogine, ``no formulation of the laws of nature that does not take into account this constructive role of time can ever be satisfactory” [75].
To address this fundamental problem, we distinguish here two different concepts of time, which we call geometric time and creative time, respectively. We show that those stem from our fundamental assumption of physics as being either deterministic or indeterministic at the fundamental, ontological level and that they both seem to contribute to our understanding of natural phenomena. In particular, we will see that geometric time is the parametric time appearing in the (deterministic) equations of motions of physics, just a coordinate that, together with the three spatial directions, labels fixed events in a geometrical block-universe.
On the other hand, creative time arises from the assumption that certain events, which are fundamentally indeterminate, become determinate. Our general approach is admittedly influenced by quantum mechanics, which as a matter of fact brought the notions of fundamental indeterminacy and indeterminism into physics. More abstractly, we consider the possibility of ontic indeterminacy: a metaphysical feature of a physical quantity that exhibits “worldly imprecision” [62, 63, 4, 85]. An ontologically indeterminate reading of the quantum state is, arguably, a widely held option (setting aside nonlocal hidden‐variable theories or many‐worlds). Although not a logical necessity, Heisenberg uncertainty together with violations of Bell inequalities provide strong reasons to regard quantum indeterminacy as more than a reasonable hypothesis. Moreover, we have shown in detail that, even at the classical level, if one adopts a finiteness‐of‐information principle, there is a consistent way to interpret classical theory as fundamentally indeterminate [46, 24, 29, 26]. Throughout the paper, we therefore take seriously the working hypothesis of ontic (sometimes called fundamental) indeterminacy.
In general, indeterminacy is a kinematic property of the state describing such a quantity. There is nothing subjectively uncertain here; in our naturalistic view, Nature simply has not yet fully determinate all objects. Indeterminism, on the other hand, refers to the possibility in Nature of having different potential (mutually exclusive) future states, given the present state and the laws of physics. Note that ontic indeterminacy in Nature is sufficient (though not necessary) for the existence of multiple potential future states, i.e., for indeterminism.
Indeterminacy can be formalized in terms of the information existing about, for example, a certain physical property. In fact, we contend that the world is composed of material entities – e.g., particles and fields – but these do not exist simpliciter. What exists is structure: information about the arrangement of things, their relations, their positions, etc. This information is objective, existing in nature independently of any observer. It is physically instantiated in the degrees of freedom of material systems: a physicalist interpretation in the spirit of Landauer’s dictum that information is physical [52].
More information can be created when a potential future is actualized, thereby reducing indeterminacy. It is the actualization of potentialities – when new information that previously did not exist is created – that makes creative time tick. This is in line with the ideas of mathematician A. Connes, who states: ”It’s quantum effervescence that generates the passage of time, not the other way around.” [22].
This idea of creative time also resonates to some extent with the positions of the philosopher H. Bergson, well known for having confronted Einstein about the nature of time, who stated: ``Time is what prevents everything from being given all at once. […] It must therefore be development. Wouldn’t it then be the vehicle of creation and choice? Would time’s existence not prove that things are indeterminate? Would time not be this very indeterminacy?” [11]. However, while vaguely inspired by Bergson, as physicists we twist the story around and adopt a more explicitly naturalistic view: it is the natural change in metaphysical indeterminacy (i.e., the creation of novel information and the passage from a potentiality to an actual property) that defines our creative time. We can thus say that, while geometric time “passes”, creative time “processes” or “occurs”, as the consequence of changes in the world. [^1] In the literature one often finds that time ``unfolds” or that time ``flows” to indicate that time is associated with something that really changes, and thus with creative time. Here we prefer to use the more active terms ``time occurs” or ``time processes”, because a film already shot does unfold, but nothing new is created there; whereas to flow is a somewhat misleading term, for it recalls the concept of speed, which already presupposes time.
@@
Figure 1: Typical evolution of a classical chaotic system in phase space. The initial state (a) has some ontic indeterminacy, i.e., the position and momentum are in this case not determinate further than the finite region depicted in figure. When evolved through the deterministic equations of motions (b), the indeterminacy spreads (while preserving the same volume), leading to multiple potential future evolutions, i.e. to indeterminism. However, only one of these potential evolutions will actualize (c), while ruling out the other previously potential futures (dashed); it is this process of actualization that makes creative time process.
@@
Starting from physics (Sect. II), we apply these two concepts of time to different disciplines, showing that this distinction gives rise to natural parallels in mathematics (classical versus intuitionistic; Sect. III), helps clarify the logic of scientific propositions (Sect. IV), and it contributes further to the metaphysics of time (so-called A-theory versus B-theory; Sect. V). More specifically, the approach based on consecutive actualizations of potentialities, that is when creative time is at work, leads to a naturalistic description of the difference between past, present, and future.
We do not claim that the account we present here constitutes a new philosophy of time, nor that it resolves enduring debates, such as that between phenomenalism and scientism, regarding our perception or the intrinsic nature of time. Rather, we aim to emphasize the distinction between two different concepts of time, which we believe arise naturally from physical science and have parallels in various other disciplines.
@@ II Two times in physics
In order to illustrate different concepts of time in physics, let us start from an example. Let us ask the following scientific question: what will be the weather like, say on the field hockey pitch of Geneva, exactly in one month from now? One can envision several procedures to reply to this question. On the one hand, it is possible to use the equations of motion of all the involved molecules of air, water vapor, etc., to calculate their evolution and so predict the weather in Geneva in a month’s time (thereby applying an algorithm, which is by definition a set of finite instructions). On the other hand, one can wait exactly one month, let the process develop, and finally observe the weather in Geneva. In-between, one could spend one week to collect more data and only then start computing the evolution of all air molecules. Let us concentrate on the two first mentioned extremal procedures. The question that we ask here is: Are these two procedures talking about the same unique concept of time, or are there multiple ones that play a role in physics? We contend that there are (at least) two different concepts of time, as exemplified by the two procedures above.
@@
Before elaborating on that, we ought to distinguish between different fundamental interpretations of physics, which become especially relevant in dealing with complex systems like the one determining the weather in our example. The standard narrative, (i) has it that classical physics is deterministic, i.e. that each state is mapped one-to-one to all other possible states at all times (with time being the parameter that appears in the equations of motion). We would like to stress that this is arguably based on a popular misconception, while, in fact, most physical theories can be regarded as indeterministic [36]. In fact, the belief that classical physics is deterministic is based on two independent ingredients: the first is that since the equations of motion are ordinary differential equations, their solutions generally have a unique continuation in the parameter time, [^2] In exceptional cases, classical equations of motion of even simple systems do not have a unique (deterministic) solution, such as in the example of Norton’s dome [66]. Moreover, in highly complex systems – such as Lagrangian fluid particles in high Reynolds-number turbulence – the fact that deterministic trajectories are not unique is rather common (see [41] and references therein). given appropriate initial conditions; the second is that the initial conditions exist fully determinate with infinite precision (i.e., a state is represented by a dimensionless mathematical point in phase space, in turn described by an n-tuple of real numbers). Only taking these two assumptions together justifies the purported fundamental determinism of classical physics. Another possibility, (ii) is to keep the idea that the system is fundamentally deterministic, but realistically it is not possible to know the initial conditions with infinite precision. This becomes most relevant in chaotic systems (like the weather in our example) in which the future behavior of the considered system highly depends on the infinitesimal digits of the initial conditions. In this case one can only statistically predict what are the possible future states compatible with the partial knowledge of the initial conditions. Finally, (iii) one can instead uphold the more radical assumption that the initial conditions (alternatively, the state of a system) simply do not exist determinate with infinite precision, but there is an ontic indeterminacy in nature (for a collection of works on this see, [33, 24, 46, 48, 29, 26]; see also [20] and references therein for recent developments on metaphysical indeterminacy). In this case, although the equations of motion may remain unchanged, the system does not undergo a necessary evolution, but multiple future states are possible (perhaps with a determinate objective tendency or propensity; see [26]), i.e., it is the ontic indeterminacy that leads to fundamental indeterminism (see Fig. 1). In what follows, when referring to indeterminacy we will have this kind of ontic property in mind and its consequent indeterminism.
@@
Let us now go back to analyze the concepts of time that are at work in the example of the weather in Geneva in a month. The first procedure, under the ideal assumption of determinism (i), merely makes use of time as a parameter that simply indicates the coordinate of a certain evolution (e.g., sunny weather in one month from the start of the computation). Together with the three spatial coordinates (i.e., in Geneva in our example), this kind of time forms a manifold that contains all the events: a four-dimensional block universe picture where past, present, and future are all on the same metaphysical footing (this is also known as Eternalism in the philosophy of time; see Sect. V). This can thus be called geometric time, the time that most physicists have come to think of as fundamental, both from classical physics and even more from relativity theory where time becomes “just” another dimension like space. [^3] Though with a different signature in the metric. In this sense, time is a relational label that allows one to position an event (as earlier or later than to another point of reference), but there is no new information created in the world.
@@
On the other hand, let us consider the case of ontic indeterminacy in the physical state (iii). There, the actual development of a non-necessary physical process – such as the evolution of a highly chaotic system like the one that determines the weather, or of a quantum system – involves, in our opinion, a time that really ``processes”. That is, a concept of time that sets an actual separation of the future from the past (with the present at their interface). And it does so when new information that was not existing before gets created – i.e., even having access to the whole existing information of the initial conditions (even of the whole state of the universe) before, would not determine all the future physical properties of a system afterwards. Again, we stress that we are not talking here about epistemic uncertainty, but rather we explicitly take a physicalist stance on information: it quantifies something objectively encoded in the degrees of freedom of the universe. In this sense, information is created when non-necessary physical events obtain, independently of any observer. Note that – as in quantum physics – even given the pure state of a system – i.e., in Butterfield’s words, “a system’s maximal (or ‘complete’) set of intrinsic (or ‘possessed’) properties” [18] – this does not fully determine the future evolution (see also Ref. [62]). Here, complete means that a state contains the maximal information presently available “in nature” about each potential property, i.e. the collection of all propensities. Yet only one potentiality, among its alternatives, eventually actualizes. The concept of time that arises from this can thus be called creative time. [^4] The distinction between geometric and creative time was sketched in Refs. [44, 25, 24, 81] and previously also in Ref. [49] where the creative time was called ``historical time”. On a similar note, G. Ellis remarked that ``The time that is the present at this instant will be the past at the next instant. […] It is this fundamental feature of time that is not encapsulated by today’s theoretical physics” [38] (see also [30]).
@@
Let us emphasize that the concept of creative time has nothing to do with language, nor with subjective or psychological processes; it ticks due to objective processes that happen by themselves in nature. Accordingly, there are two types of events: necessary events (e.g., when two geometric worldlines cross each other) and non-necessary events which occur when potentialities become actual. This is exemplified, for instance, by the case of a single photon impinging on an array of detectors—eventually, only one detector clicks.
@@
So, coming back to the example of the weather in Geneva in a month, one can carry out either of the envisioned procedures to determine the outcome – (i) compute the state of the weather in a month using the relevant laws of physics, or (ii) wait and directly observe the weather. Both these procedures take some time. In practice, to carry out the first procedure, one maps the time interval that it takes to perform the computation $\Delta t^{\prime}$ to the time appearing in the equations of motion. That is, to the interval $\Delta t=t_{f}-t_{i}$, where $t_{i}$ is the parametric time associated to the initial state and $t_{f}$ is the parametric time of the considered final state (i.e., the moment when we want to know the weather in Geneva). Note that both $\Delta t^{\prime}$ and $\Delta t$ are geometric times. For this to be a prediction, the (deterministic) computation should last $\Delta t^{\prime}<\Delta t=1$ month, and return as an outcome the weather in Geneva at $t_{f}$, say ``sunny”. [^5] In principle the time for the “prediction” could take longer than the time interval to observe the outcome. In such a case, one would anyways ask for consistency of the calculation afterwards.
@@
One can also think of more subtle questions about the weather, such as ``will there be (anytime in history) 700 consecutive days of rain in Geneva?”. If the complex system that forms the weather is fully deterministic, answering this question seems only a matter of geometric time, although the answer may remain undecided for any finite value of geometric time. One can run the computation of the dynamics of the system and at some point find that the answer is ``yes”, if one finds the 700 consecutive days of rain, otherwise the algorithm will keep searching without ever halting (however, this question might be more complex, see Sect. III).
@@
Note that $\Delta t^{\prime}$ can be sped up by a more efficient computation, namely, this depends on how fast one ``runs” through the parameter time as it appears in the equations of motion (this depends on the used algorithm and, physically, one can really think of the speed of the head of a Turing machine in reading the tape). Moreover, one can use any other effects that allow one to manipulate geometric time, such as time dilation in relativity that modifies the interval of geometric time $\Delta t^{\prime}$ (see Subsect. VI.1). So, the map $\Delta t\rightarrow\Delta t^{\prime}$ is a one-to-many function.
@@
In contrast, in the process involving creative time, then one has no choice but to wait a month and see what happens, because new information is created along the way through a series of indeterministic processes. So, the initial and the final states are not deterministically mapped to each other. Creative time is required when an actualization event of the values of physical variables that were previously indeterminate occurs. It is the process of change from indeterminate to determinate – with an associated creation of novel information – that brings about this “occurrence” of time: creative time ticks when new information is created.